OpenStudy (anonymous):

what is the value of the discriminate -18x^2+5x-13=0 I have been doing this problem and I got to what I thought may be the answer but it just seems to outragious to be true

OpenStudy (anonymous):

okay discriminant is b^2 - 4ac, so in this case b = 5, a = -18, c = -13, so 5² - 4*-18*-13 = 25-936=-911

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

see I came out with -982.12? I used -b/2(a) so I did -5/2(-18)= -5/-36

OpenStudy (anonymous):

if you got anegative answer in the discriminant it has a complex roots

OpenStudy (anonymous):

b/2a is the vertex or the min or maximum value of the quadratic equation

OpenStudy (anonymous):

ok we use this and then take the answer and put it in the equation is what I thought

OpenStudy (anonymous):

the vertex x=-b/(2a)=-5/(2)(-18)=5/36

OpenStudy (anonymous):

you can plug this to the orig eq y=-18x^2+5x-13 y=-18(5/36)^2 +5(5/36)-13

OpenStudy (anonymous):

y=-287/24 =-11.9583333

OpenStudy (anonymous):

s0 your (x,y)=(5/36, -287/24)

OpenStudy (anonymous):

s0 your (x,y)=(5/36, -287/24)

OpenStudy (anonymous):

thats the vertex o max value of the equation y=-18x^2+5x-13

OpenStudy (anonymous):

but the prob is asking for the discriminate

OpenStudy (anonymous):

b^2 - 4ac=-911